OpenStudy (starryknight):

Identify the prime number.(metal for answer) 39 33 13 25

OpenStudy (anonymous):

You know what a prime number is?

OpenStudy (bookworm14):

A prime number is a number that can only be multiplied by 1 and itself 3*13=39 (composite) 5*5=25 (composite) 3*11=33 (composite) 1*13= 13 (prime) Hope this helps !!- Bree :)

OpenStudy (anonymous):

A Prime Number can be divided evenly only by 1, or itself. And it must be a whole number greater than 1. This is the right definition, Bookworm14.

OpenStudy (bookworm14):

I know the definition, I just gave a simple version that was easy to understand and memorize, still just as effective , thank you very much :) @d3Xter